0-07-865113-1
algebra equation 48(3 1/6)
How will break this down?

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(21584)

algebra equation 48(3 1/6)
How will break this down?
-----------
All equations have an equal sign. I don't know what you're trying to do here.

Answer by stanbon(48545)

48(3 1/6)
Multiply to get 48 * (19/6) = 8*19 = 152
-----------------------
Cheers,
Stan H.